OpenStudy (anonymous):

Use the quadratic equation y = x^2 + 3x – 10. Where does the vertex of the parabola lie? A. below the x-axis B. above the x-axis C. on the x-axis D. on the y-axis

OpenStudy (anonymous):

first coordinate of the vertex is always \(-\frac{b}{2a}\) in your case it is \[-\frac{3}{2}\] to find the second coordinate, replace \(x\) by \(-\frac{3}{2}\) and see what you get

OpenStudy (anonymous):

i am fairly sure that the answer will be negative, so it lies below the x - axis

OpenStudy (anonymous):

ty